PROPOSAL TO

DEPARTMENT OF ENERGY

NATIONAL NUCLEAR SECURITY ADMINISTRATION

ADVANCED SIMULATION AND COMPUTING PROGRAM

ACADEMIC STRATEGIC ALLIANCE PROGRAM

FOR

TWO-YEAR RAMP-DOWN FUNDING

OF THE

CENTER FOR ASTROPHYSICAL THERMONUCLEAR FLASHES

Submitted by The University of Chicago 5801 South Ellis Avenue Chicago, IL 60637

For the period 1 April 2008 through 30 September 2010

February 2008

Amount: $3,000,000

Principal Investigator:

Professor Don Q. Lamb Enrico Fermi Institute and Dept. Astronomy & Astrophysics

(312) 702-7194

Carol Zuiches Associate Vice-Pres. for Res. & Director University Research Administration

(312) 702-8604